A sample of hydrogen gas is kept in a container fitted with a movable piston set at the 8.0-liter mark at 50 degrees Celsius and


110 kPa. At what temperature in Kelvin will the piston to move to the 9.0-liter mark without a change is the pressure reading?

1
Expert's answer
2021-11-12T13:34:02-0500

"V_1 = 8.0 \; L \\

T_1 = 50 + 273 = 323 \; K \\

p = 110 \;kPa = constant \\

V_2 = 9.0 \; L \\

\frac{p_1V_1}{T_1} = \frac{p_2V_2}{T_2} \\

\frac{V_1}{T_1} = \frac{V_2}{T_2} \\

T_2 = \frac{V_2T_1}{V_1} \\

T_2 = \frac{9.0 \times 323}{8.0} = 363.3 \;K"

Answer: 363.3 K